Could someone please help me with my algebra? Here is my question I'm not sure how to set it up either.
Find the selling price per pound of a tea mixture made from 40lb of tea costing $5.40 per pound and 60lb of tea costing $3.25 per pound. Thanks for your help.

Answer by stanbon(57347)

Find the selling price per pound of a tea mixture made from 40lb of tea costing $5.40 per pound and 60lb of tea costing $3.25 per pound
-------------
$5.40 tea DATA:
Amount = 40 lbs. ; Value = 5.40*40 = 216 dollars
---------------
$3.25 tea DATA:
Amount = 60 lbs. ; Value = 3.25*60 = 195 dollars
---------------
Mixture DATA at x dollars per pound
Amount = 100 lbs ; value = 100x dollars
===================
EQUATION:
value + value = value
216 + 195 = 100x
411 = 100x
x = $4.11 per pound
==========================
Cheers,
Stan H.